Accelerator Update for December 12, 2008 - December 15, 2008
The Integrated Luminosity for 12/8/08 to 12/15/08 was 73.73 Pb-1. This set a new record for integrated luminosity.
Friday Dec 12
The day shift began with the Tevatron (TeV) in shot setup, with Operations monitoring the Recycler stash (9.5E10), with the Antiproton Source stacking (65.0E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments waiting for the next store, and with the MiniBooNE experiment taking beam. NuMI was getting ready to operate with their new horn.
Operations established store 6634 at 8:22 AM with an initial luminosity of 335.87E30.
At 10:18 AM, the P1 transfer line beam permit tripped off due to a power supply’s current (I:HT701) drifting too high. Operators reset the permit and loaded a new file for the supply.
The Neutron Therapy Facility reported at 10:39 AM that they had completed treatment on their last patient for the day.
Operators reported at 12:20 PM that they had searched and secured NuMI’s MI-65 enclosure. NuMI had a beam permit four minutes later.
Operations established beam to NuMI at 12:36 PM. NuMI experts began conducting target scans.
Operations reported at 3:58 PM that the Recycler had stashed antiprotons seven times during the day shift.
Operations disabled beam to MTest at 6 PM.
Operations terminated store 6634 at 11:38 PM.
Operations put the TeV into shot setup at 11:52 PM.
Operations reported at 11:57 PM that the Recycler had stashed eight times during the evening shift.
Saturday Dec 13
The midnight shift began with the TeV in shot setup, with Operations monitoring the stash (372.2E10), with Pbar stacking (35.9E10), and with MiniBooNE and NuMI taking beam.
Operations established store 6635 at 1:25 AM with an initial luminosity of 337.8E30.
Operations reported at 8:03 AM that the Recycler had stashed seven times during the midnight shift.
Operations established beam to MTest at 8:17 AM.
Operations put Pbar into standby at 12:34 PM.
Operations increased the current of the remaining Booster kickers (to make up for the missing kicker) at 12:49 PM and resumed sending beam to Pbar, NuMI, and MiniBooNE.
Booster experts reported at 1:47 PM that MKS08 had a shorted power supply transformer cable. They began repairs.
Operations terminated store 6635 at 2:06 PM.
Operations put Pbar into standby at 2:11 PM for the Booster kicker work.
Operations prepared D0 and CDF for access at 2:21 PM.
Booster resumed sending beam at 2:44 PM. Booster experts reported that they also worked on MKS07.
At 3:33 PM, a MI RF station (MIRF14) tripped off for the first of four times in about 30 minutes. Several of the trips took down the NuMI beam permit. Everything reset.
CDF and D0 were out of access by 4:22 PM.
Operations put the TeV through a dry squeeze at 5:37 PM.
Operations disabled beam to MTest around 6 PM.
Operations put the TeV through a wet squeeze at 6:15 PM.
Operations checked the TeV coalescing at 6:56 PM.
Operations put the TeV through another wet squeeze and checked the orbits at 6:57 PM.
Operations put the TeV into shot setup at 7:18 PM.
Operations reported at 7:58 PM that the Recycler had stashed nine times during the last 12 hours.
Operations established store 6639 at 9:41 PM with an initial luminosity of 348.2E30.
Sunday Dec 14
The midnight shift began with Operations monitoring store 6639 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.
At 2:13 AM, a Linac Klystron RF station (KRF3) tripped off. It reset.
Operations reported at 7:58 AM that the Recycler had stashed 11 times during the last 12 hours.
Operations established beam to MTest at 8:40 AM.
Operations terminated store 6639 at 12:46 PM.
Operations put the TeV into shot setup at 12:59 PM.
At 2:18 PM, the Booster kickers MKS07 and 08 tripped off for the first of two times in four minutes. Operators talked to an expert, located a spare filament power supply, and replaced the existing supply for the kickers.
Booster resumed sending beam at 2:46 PM.
At 3:04 PM, Operations reported that the Booster kicker power supplies (MKS07 and 08) were not regulating properly. This caused several failed attempts to load protons for the next store.
Operations established store 6640 at 3:50 PM with an initial luminosity of 316.77E30.
Operations disabled the MTest beam at 6:18 PM.
At 7:35 PM, Operations reported that the Pelletron solenoids were not taking settings. Operations contacted experts. This put a halt to stashing antiprotons. It was determined that this was a Front End problem. Operations contacted a controls expert.
Operations reported at 8:07 PM that the Recycler had stashed eight times during the last 12 hours.
At 9:56 PM, KRF3 tripped off. It reset.
Monday Dec 15
The midnight shift began with Operations monitoring store 6640 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.
At 12:39 AM, KRF3 tripped off. It reset.
Operations terminated store 6640 at 6:53 AM.
Operations put the TeV into shot setup at 7:04 AM.
Operations reported at 8:09 AM that the Recycler had stashed 11 times during the last 12 hours.
The Plans for Monday
The plans for today are to stack, stash, and store. The Run Coordinator will allow Booster experts time to further investigate their kicker problems and for Pbar experts to conduct a Pbar Debuncher cooling study.
Comments and Suggestions
|last modified 01/05/2009 email Fermilab|
|Security, Privacy, Legal|